Find the critical values for the following confidence intervals (Z-scores)    (Please show your work)

a.) 99% b.) 98% c.) 90%
Answer #1

solution

a.

At 99% critical value  is ,

  = 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

Z/2 = Z0.005 = 2.576 ( Using z table )

b.

At 98%critical value  is ,

= 1 - 98% = 1 - 0.98 = 0.02

/ 2 = 0.02/ 2 = 0.01

Z/2 = Z0.01 = 2.326 ( Using z table    )

c.

At 90% critical value  is   

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645 ( Using z table )